x+2y-z=-7
2x-2y-z=6
x+y-2z=-6

Answer by stanbon(75887)

x+2y-z=-7
2x-2y-z=6
x+y-2z=-6
------------------
Use any method you know to get:
x = 1
y = -3
z = 2
=========
Cheers,
Stan H.